OpenStudy (anonymous):

Find the general solution to the equation: u'' - u = 4 - sin(t) -2e^t

OpenStudy (anonymous):

Okay, firstly you want to solve the homogeneous part of the equation (i.e, where u''-u=0). So do that: \[\lambda^2-1=0 \rightarrow \lambda = \pm 1\] So your homogeneous solution is: \[u_H(t)=c_1e^{-t}+c_2e^{t}\] Then for the second part use undetermined coefficients. You want to look for something in the form: \[y_P(t)=A+B \sin(t)+D \cos(t)+E t e^t\] NOTE: The t on the e^t comes from the fact that there is an e^t in your homogeneous solution.

OpenStudy (anonymous):

From there, you want to differentiation your particular solution twice and then plug it in to solve for you coefficients. \[y'_P(t)=B \cos(t)-D \sin(t)+E(e^t+te^t)\] \[y''_P(t)=-B \sin(t)-D \cos(t)+E(e^t+e^t+te^t)\] Plug this in and solve for the coefficients: \[-B \sin(t)-D \cos(t)+E(e^t+e^t+te^t)-A-B \sin(t)-D \cos(t)-Ete^t=\]\[4-\sin(t)-2e^t\] From here group all your terms together using facts from linear algebra so you know that: -B-B=-1 -D-D=0 E+E=-2 -A=4 So your general solution is the sum of the homogeneous and the particular giving: \[c_1e^t+c_2e^{-t}-4+\frac{1}{2}\sin(t)-te^t\] Here is the wolfram page to confirm it: http://www.wolframalpha.com/input/?i=solve+y''-y%3D4-sin(t)-2e^t